OpenStudy (anonymous):

Please help me to simplify without a negative exponent: (b^3 c^1/4)^-4/5

OpenStudy (anonymous):

what do you mean? are you solving for a specific variable or anything?

OpenStudy (anonymous):

I am simplifying a quadratic formula. If I multiply ^-1/4 by the exponents in the parentheses I got: 1/5b^12c

OpenStudy (anonymous):

\[(b^{3}c ^{1/4})^{-4/5}\] im sorry i still dont follow what you are asking there is a couple ways you can manipulate it i believe

OpenStudy (anonymous):

The question is asking me to simplify a quadratic expression by multiplying and to solve without using negative exponents in the answer. (b^5/3 c^-1/2)^3

OpenStudy (anonymous):

\[a^{-n} = \frac{1}{a^n}\] \[\frac{1}{a^{-n}}= a^n\]

OpenStudy (anonymous):

\[(b^{3} c^{1/4})^{-4/5}\] Multiply all the exponents inside by the -4/5: \[b^{-12/5} c^{-1/5}\] To eliminate the negative exponents move both terms into the denominator: \[\frac{ 1 }{ b^{12/5} c^{1/5} }\]

OpenStudy (anonymous):

I was close! Here is your medal.